3) You are asked for a loan and promised repayment of $250 two years from now. If the minimum interest rate you will accept is 8%, what is the maximum amount you will loan? a.) $119 b.) $126 c.)$129 d.) $139 e.)$150
Answer #1

e). $150

First we need to find out the present value of future cash flow $250., using formula;

present value = future value/(1+ r)n

where , r = interest rate   

. n = number of periods

PV = 250/(1+0.08)2 = 250/(1.08)2 = 250/1.1664 = 214.334

Since the present value of $250 at 8% after two years is 214.334 . This is maximum value he can loan if repament value is $250 after two years. But among the options , the values are less than 214.334. So the highest value among the options I.e. $150 will give as loan.

​​​​​​